OpenStudy (anonymous):

a(1)=7 an=a(n-1) -4 how do i find the first 5 numbers in this sequence?

OpenStudy (mathstudent55):

What does this rule mean? \(\large a_n=a_{n-1} -4\)

OpenStudy (anonymous):

I dont know, I was also given that n is greater than or equal to 2

OpenStudy (mathstudent55):

Ok. This is how you interpret the rule. \(\large a_n\) means any therm in the sequence. \(\large a_{n - 1} \) means the term before \(\large a_n\) That means that any term in the sequence is the previous term minus 4.

OpenStudy (anonymous):

so 3 is 2-4? and 4 is 3-4? and 5 is 5-4?

OpenStudy (mathstudent55):

The first term is given to you. \(\large a_1 = 7\) Now we need to find the second term. The second term is the previous term minus 4. \(\large a_2 = a_1 - 4 = 7 - 4 = 3\)

OpenStudy (mathstudent55):

n is not the value of the term. n is simply a way to keep count of each term.

OpenStudy (anonymous):

ohhh, so then